OpenStudy (luigi0210):

Evaluate:

OpenStudy (luigi0210):

\[\int\limits (2-3e^x)~dx \]

OpenStudy (yttrium):

Is it indefinite integral?

OpenStudy (luigi0210):

Yup, another simple one.

OpenStudy (yttrium):

Just distribute the dx and integrate it seperately. I think you can do this! :)

OpenStudy (yttrium):

Do I need to say the answer?

OpenStudy (luigi0210):

\[\large \int\limits 2dx-3~\int\limits e^x=2x-3e^x+C\]

OpenStudy (yttrium):

Good! You got it correctly! :)

OpenStudy (luigi0210):

I'm just bored. ._.

OpenStudy (yttrium):

LOL HAHAHA

OpenStudy (luigi0210):

Oh well.